### v1

@@ 2.4 Global Explanations of Supervised Learning with Shapley Values
So far, we have used Shapley values to provide local explanations, which describe how individual feature values contribute to a single prediction $f(x)$. However, one may also want to understand how feature values influence predictions in general, leading to what are known as global explanations. While global Shapley-based explanations of model predictions have not been explored in the literature, similar methods have been used to explain a model’s loss.
@@
$v^{\ell}(\mathcal{C})=\mathbb{E}_{p(x)}\left[v^{\ell}_{x}(\mathcal{C})\right],$
@@
$\phi_{i}(v^{\ell})=\mathbb{E}_{p(x)}\left[\phi_{i}(v^{\ell}_{x})\right]$
@@ 2.5 Computing Shapley Values in Practice
where $\mathcal{C}$ is the set of features that precede $i$ in ordering $O$ and $O$ is sampled uniformly from $\pi({\cal F})$, the set of all permutations of ${\cal F}$. The second method is the kernel method, which approximates characteristic values using Equation 3, rewrites the Shapley values of Equation 1 as the solution to a particular weighted least squares problem, and then solves for the Shapley values for all features at a point $x$ simultaneously. The kernel method was introduced by Lundberg and Lee (2017); we direct the reader to Aas et al. (2021) for a detailed description of the approach.
@@
3 Explaining Reinforcement Learning
A fundamental question in explaining reinforcement learning is: What exactly should be explained? Here, we propose one possible answer—explaining an agent’s interaction with its environment requires explaining three key elements:
  1. 1.
    Behaviour: How does the agent choose actions?
  2. 2.
    Performance: What reward do these actions earn?
  3. 3.
    Value estimation: How can we estimate these rewards?
@@ 3 Explaining Reinforcement Learning
Each of these elements provides a different perspective on the reinforcement learning problem. Together, they help us build a comprehensive understanding of an agent’s behaviour, the consequences of its decisions, and our ability to anticipate those consequences.
Consider an autonomous vehicle navigating a city using road signs (see Figure 1). The signs provide directional guidance and distance information, influencing the vehicle’s decision-making process. By analysing how the agent interprets these cues, selects actions, and achieves its goal, we can examine reinforcement learning through the following three perspectives.
@@
Figure 1: An autonomous vehicle navigates toward a destination marked by flags, following road signs that provide directions and distances. The vehicle’s policy follows the black path, turning right at the first junction, while the red path represents an alternative shortest route.
  • Explaining Behaviour. An agent’s behaviour describes how it selects actions given its current knowledge. In our example (see Figure 1), the autonomous vehicle follows a policy that dictates its route based on road signs—such as turning right at the first junction because a sign indicates that direction. The vehicle might take a different route if a sign is missing or incorrect. Explaining behaviour means understanding why an agent makes certain decisions based on the available information.
  • Explaining Performance. Performance measures how well the agent achieves its objective, typically evaluated through the expected return of its policy. For instance, if the autonomous vehicle reaches its destination in minimal time, we consider it to have high performance. Performance is distinct from behaviour: a change in behaviour does not necessarily result in a change in performance. In Figure 1, the vehicle can turn either right or left at the first junction, but both paths lead to the destination in the same distance. While behaviour explanation concerns how an agent selects actions, performance explanation concerns how well those actions achieve the goal.
  • Explaining Value Prediction. Value estimation involves estimating how successful a behaviour will be. Consider the road sign distances in Figure 1. The first sign indicates that the destination is 10 miles away, while the second sign shows it is only 2 miles away. We expect a higher expected return at 2 miles than at 10 miles because the agent is closer to its goal. Importantly, estimating value considers how different inputs affect our estimate of expected return without necessarily changing the agent’s actual behaviour—the policy itself does not rely on the road sign distances. This distinguishes value estimation from behaviour and performance: the latter describes what actually happens, while the former describes what we expect to happen.
@@
By explaining behaviour, performance, and value estimates, we gain a holistic understanding of reinforcement learning—covering an agent’s behaviour, the returns of that behaviour, and how we estimate those returns. In the following sections, we formalise these explanations using Shapley values. Specifically, we analyse how to explain behaviour (Section 4), performance (Section 5), and value estimation (Section 6), introducing our framework: Shapley Values for Explaining Reinforcement Learning (SVERL).
@@ 4.1 State Features Collaborate to Determine Action Probabilities
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\tilde{\pi}_{s}^{a}(\mathcal{C})$ returns the probability of selecting action $a$ in state $s$ when the features $\mathcal{C}\subseteq{\cal F}$ are known by the agent and the features $\bar{\mathcal{C}}={\cal F}\setminus\mathcal{C}$ are unknown.
@@
Central to the definition of the discrete policy game $({\cal F},\tilde{\pi}_{s}^{a})$ is the characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow[0,1]$, which specifies how the agent would act—that is, the probability of selecting action $a$ when in state $s$—when only a subset of the feature values is known. This function depends on the underlying policy $\pi$, which governs behaviour when all state features are fully observed. However, the policy $\pi$ does not define the agent’s behaviour when some feature values are unknown. To account for this, we introduce the characteristic function $\tilde{\pi}_{s}^{a}$, which extends $\pi$ to settings where the agent has partial knowledge of the state. We use the same symbol $\pi$ to highlight this relationship: in the case where all features are known, that is, $\mathcal{C}={\cal F}$, we require that $\tilde{\pi}_{s}^{a}(\mathcal{C})=\pi(s,a)$. The precise construction of the characteristic function $\tilde{\pi}_{s}^{a}$ will be addressed later in Section 4.2. For now, we assume it is given and turn to the central question of allocating credit among the features.
@@
The question we now face is: What is the precise contribution of each state feature value $s_{i}\in\mathcal{S}_{i}$ to the probability of selecting action $a$ in state $s$? We denote this contribution by $\phi_{i}(\tilde{\pi}_{s}^{a})$. In other words, how can we fairly distribute—across the state features—the difference between the action probability when all feature values are known and the action probability when none are known? Below, we define four axioms that any attribution $\phi_{i}(\tilde{\pi}_{s}^{a})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation. When referring to a feature’s contribution, we implicitly mean the contribution of its value $s_{i}$ at state $s$.
@@
  • Axiom 1 (Efficiency). The difference in value between the action probability when all features are known, $\tilde{\pi}_{s}^{a}({\cal F})$, and the action probability when no features are known, $\tilde{\pi}_{s}^{a}(\emptyset)$, is fully distributed over the features ${\cal F}$:
    $\tilde{\pi}_{s}^{a}({\cal F})-\tilde{\pi}_{s}^{a}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\pi}_{s}^{a})$
@@
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi+(1-\alpha)\,\omega$ that follows, at each decision stage, policy $\pi$ with probability $\alpha$ and policy $\omega$ with probability $(1-\alpha)$, a state feature’s contribution to the action probability $\alpha\,\pi(s,a)+(1-\alpha)\,\omega(s,a)$ is equal to the weighted sum of its contributions to the individual action probabilities $\pi(s,a)$ and $\omega(s,a)$:
@@
  • Axiom 3 (Symmetry). Two features are assigned the same contribution to the action probability $\pi(s,a)$ if they make equal marginal contributions to the probability of selecting action $a$ in state $s$ for all possible coalitions of features:
@@
  • Axiom 4 (Nullity). A feature is assigned zero contribution to the action probability $\pi(s,a)$ if it makes zero marginal contribution to the probability of selecting action $a$ in state $s$ for all possible coalitions of feature values:
@@
These four axioms are directly analogous to the standard axioms from cooperative game theory used to fairly assign player contributions to the outcome of a coalitional game, as discussed in Section 2.
@@
We now revisit the example introduced in Sections 3 and 1, where an autonomous vehicle navigates through junctions using road signs that indicate both the direction and distance to the destination. At the first junction, the vehicle turns right. How can we explain this behaviour? We apply our framework to attribute the decision to the observed features at that state. However, we must first consider how the agent might behave when one or more features are unknown.
To explore this, we look at two possible ways the agent could act at the first junction when features are unknown. For clarity, we refer to these behaviours as separate agents. Agent A bases its decisions solely on the direction the sign indicates; if the direction is visible, it follows it. Otherwise, it chooses randomly. Agent B uses both direction and distance to determine which way to turn; if the direction or distance is visible in the first state, it moves right. Otherwise, it also chooses randomly. These assumptions give rise to different policies under unknown information, shown in Table 1.
Table 1: Agent behaviour under different subsets of observed features. Each cell shows the probability of turning left (L) and right (R).
@@
A $\pi(\text{L})=0,\ \pi(\text{R})=1$ 0, 1010,\ 10 , 1 0.5, $,\ 0.5$ 0.5 0.5, $,\ 0.5$ 0.5
B 0, 1010,\ 10 , 1 0, 1010,\ 10 , 1 0, 1010,\ 10 , 1 0.5, $,\ 0.5$ 0.5
We then compute the Shapley values for each policy, attributing the agent’s choice of action to the direction and distance features. The results are shown in Table 2. For Agent A, all credit is assigned to the direction feature since the policy disregards distance entirely. For Agent B, both features contribute equally because they both support the decision to turn right. In both cases, the features that increase the probability of turning right also decrease the probability of turning left—resulting in negative Shapley values. This example illustrates how the explanation depends critically on what is assumed about agent behaviour under partial information. We next discuss how we might approach this problem in general.
Table 2: Shapley values for each feature under the two agent policies.
@@
A L $-0.5$ 00
R $0.5=\tfrac{1}{2}[(1-0.5)+(1-0.5)]$ 00
B L $-0.25$ $-0.25$
R $0.25$ $0.25$
@@ 4.2 What Can Be Assumed About Agent Behaviour When Some State Features Are Unknown?
We now consider how an agent behaves when some state features are unknown—a question that must be addressed to compute Shapley values. The agent’s policy $\pi$ defines behaviour only when all feature values are known; it does not specify how the agent acts under partial information. However, to evaluate the characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$ for all subsets $\mathcal{C}\subseteq{\cal F}$, we must make assumptions about behaviour in these cases. In what follows, we introduce one such approach and then discuss alternative modelling choices and their implications.
@@
Definition 5 (Policy Characteristic Function)
@@
$\tilde{\pi}_{s}^{a}(\mathcal{C})\equiv\mu(s_{\mathcal{C}},a)=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\pi(s^{\prime},a).$
@@
The expectation is taken over the marginal steady-state distribution $p^{\pi}(S)$—the probability that an agent following policy $\pi$ is in state $s$.
When only the features in $\mathcal{C}$ are known, the policy $\mu$ provides the closest approximation to the original policy $\pi$. As such, the difference between $\mu(s_{\mathcal{C}},a)$ and $\pi(s,a)$ quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are to reproduce the behaviour of policy $\pi$; this reflects the influence of those unknown features. Defining the characteristic function as $\tilde{\pi}_{s}^{a}(\mathcal{C})=\mu(s_{\mathcal{C}},a)$ thus ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the policy game. Because this construction relies only on the agent’s policy, not on how it was derived, it is general to any policy—optimal and non-optimal.
4.3 How Else Could Agent Behaviour Be Represented Under Partial Information?
In the previous subsection, we defined behaviour with unknown features by marginalising the policy over the conditional steady-state distribution. Here, we propose and briefly discuss several alternative approaches.
One approach is to learn policies for each subset of observed features. However, policies learnt with partial information may fail to converge or meaningfully reflect the original policy we aim to explain. Another approach is to treat “unknown” as a valid feature value, but this would require modifying the MDP so the agent can learn to interpret the “unknown” feature appropriately. Similarly, unknown features could be imputed with a placeholder, but without adapting training, the agent has no reason to interpret such values as indicating missingness. These approaches all require changes to the MDP or training process, making them restrictive and unsuitable for post hoc explanation—though they may offer promising directions for future work.
Another option is to define the characteristic function by marginalising over the joint distribution of feature values. Unlike the approaches above, this does not require changing the MDP or training process. Formally, it is defined as:
@@ 4.3 How Else Could Agent Behaviour Be Represented Under Partial Information?
This approach differs from marginalising over the conditional steady-state distribution in much the same way as in supervised learning (Section 2.3). It assumes feature independence and yields equivalent results when that assumption holds. As a result, it assigns zero Shapley value to any feature not explicitly used by the policy. This outcome may appear desirable, but it is less meaningful in many machine learning models, where all features and interdependencies are explicitly used. When features are dependent, joint marginalisation may produce combinations of values that do not correspond to any valid state. Since Shapley value attribution relies on measuring behavioural changes in response to changes in known feature values, such undefined policy queries make these changes uninterpretable. For this reason, we focus on the conditional approach in this work, although adapting our framework for joint marginalisation is straightforward. The choice of characteristic function may ultimately depend on the user, who must understand the underlying assumptions to interpret the resulting Shapley values correctly.

### v2

@@ 2.4 Global Explanations of Supervised Learning with Shapley Values
So far, we have described using Shapley values to provide local explanations, which describe how individual feature values contribute to a single prediction $f(x)$. One may also want to understand how feature values influence predictions in general, leading to what are known as global explanations. While global Shapley-based explanations of model predictions have not been explored in the literature, similar methods have been used to explain a model’s loss.
@@
$v^{\ell}(\mathcal{C})=\mathbb{E}_{p(x)}\left[\ell\left(f_{x}(\mathcal{C}),y\right)\right],$
@@
$\phi_{i}(v^{\ell})=\mathbb{E}_{p(x)}\left[\phi_{i}(v^{\ell}_{x})\right].$
@@ 2.5 Computing Shapley Values in Practice
where $\mathcal{C}$ is the set of features that precede $i$ in ordering $O$ and $O$ is sampled uniformly from $\pi({\cal F})$, the set of all permutations of ${\cal F}$. The second method is the kernel method, which approximates characteristic values using Equation 3, rewrites the Shapley values of Equation 1 as the solution to a particular weighted least squares problem, and then uses the approximate characteristic function to solve for the Shapley values of all features at a point $x$ simultaneously. The kernel method was introduced by Lundberg and Lee (2017); we direct the reader to Aas et al. (2021) for a detailed description of the approach.
@@
3 Explaining Reinforcement Learning Agents
We start our exploration by asking what would be useful to explain regarding reinforcement learning agents for people who develop, test, and deploy such agents. We propose three explanatory targets: behaviour, outcomes, and predictions.
Behaviour refers to how the agent acts, that is, the decision to take a particular action in a given situation.
Outcomes refer to measurable consequences of an agent’s behaviour, such as the next state, immediate reward, or expected return. Here we focus on expected return as the outcome of interest. The explanation framework we develop can be extended to other outcomes.
Predictions refer to estimates of outcomes. These predictions may be made by the agent itself or by an outside observer, by using information which may or may not match the information used by the agent when deciding how to act.
@@ 3 Explaining Reinforcement Learning Agents
Each of these explanatory targets—behaviour, outcomes, and predictions—provides a useful perspective on the agent. Together, they help build a comprehensive understanding of an agent’s decisions, the consequences of those decisions, and estimates of those consequences.
Consider an illustrative example depicted in Figure 1. An autonomous vehicle is navigating a city using road signs that display, at each junction, the direction and distance to the vehicle’s destination. The vehicle follows the directions shown on the road signs. For example, it turns right at the first junction. Consequently, the agent takes the path shown in black ink. An alternative shortest route is shown in red ink.
@@
Figure 1: An autonomous vehicle navigates toward a destination marked by flags. The agent takes the path shown in black ink, turning right at the first junction. The path shown in red ink represents an alternative shortest route.
Explaining behaviour means understanding and describing why an agent acts the way it does, for example, why the vehicle turns right at the first junction.
Explaining outcomes means explaining direct, measurable consequences of an agent’s behaviour. For instance, in our example, we may want to explain the total distance travelled by the vehicle. Outcomes are distinct from behaviour, and a change in behaviour does not necessarily result in a change in outcomes. For instance, in our example, at the first junction, turning right or left both lead to the destination via routes of equal length. Consequently, while the road sign at the first junction may influence how the vehicle acts, it does not influence the distance travelled by the vehicle to reach its destination.
Explaining predictions means explaining an estimate of an outcome. This differs from explaining outcomes themselves in a subtle but important way. In explaining outcomes, we explain the indirect influence of information on an outcome through its influence on behaviour. For example, different information may cause the agent to act differently, which may then lead the agent to collect a different return. In contrast, in explaining predictions, we explain the direct influence information has on predictions of outcomes, with agent behaviour not impacted. In our example, sign directions guide the agent’s decisions and thus shape its outcomes, while sign distances—though not used by the agent—can inform external predictions of those outcomes, such as predicting the agent’s expected return based on its distance from the goal. The distinction between explaining outcomes and explaining predictions will become clear throughout the paper and is illustrated in detail in Section 6.4.
@@
By explaining behaviour, outcomes, and predictions, we gain a holistic understanding of a reinforcement learning agent: how it behaves, what its behaviour produces, and how we estimate those results. In the following sections, we formalise these explanations using Shapley values. Specifically, we analyse how to explain behaviour (Section 4), outcomes (Section 5), and predictions (Section 6), introducing a unified framework: Shapley Values for Explaining Reinforcement Learning (SVERL).
@@ 4.1 State Features Collaborate to Determine Action Probabilities
A set =${\cal F}=\{1,\dots,n\}$ of features and a characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow\mathbb{R}$, where $\tilde{\pi}_{s}^{a}(\mathcal{C})$ returns the probability of selecting action $a$ in state $s$ when the values of features $\mathcal{C}\subseteq{\cal F}$ are known by the agent and the features $\bar{\mathcal{C}}={\cal F}\setminus\mathcal{C}$ are unknown.
@@
Central to the definition of the discrete policy game $({\cal F},\tilde{\pi}_{s}^{a})$ is the characteristic function $\tilde{\pi}_{s}^{a}:2^{|{\cal F}|}\rightarrow[0,1]$, which specifies how the agent would act—that is, the probability of selecting action $a$ when in state $s$—when only a subset of the feature values is known. This function depends on the underlying policy $\pi$, which governs behaviour when all state features are fully observed. However, the policy $\pi$ does not define the agent’s behaviour when some feature values are unknown. To account for this, we introduce the characteristic function $\tilde{\pi}_{s}^{a}$ to extend $\pi$ to settings where the agent has partial knowledge of the state. We use the same symbol $\pi$ to highlight this relationship: in the case where all features are known, that is, $\mathcal{C}={\cal F}$, we require that $\tilde{\pi}_{s}^{a}(\mathcal{C})=\pi(s,a)$. The precise construction of the characteristic function $\tilde{\pi}_{s}^{a}$ will be addressed later in Section 4.2. For now, we assume it is given and turn to the central question of allocating credit among the features.
@@
The question we now face is: What is the precise contribution of each state feature value $s_{i}\in\mathcal{S}_{i}$ to the probability of selecting action $a$ in state $s$? We denote this contribution by $\phi_{i}(\tilde{\pi}_{s}^{a})$. In other words, how can we fairly distribute—across the features—the difference between the action probability when all feature values are known and the action probability when none are known? Below, we define four axioms that any attribution $\phi_{i}(\tilde{\pi}_{s}^{a})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation. When referring to a feature’s contribution, we implicitly mean the contribution of its value $s_{i}$ at state $s$.
Axiom 10 (Efficiency)
@@
The difference in value between the action probability when all features are known, $\tilde{\pi}_{s}^{a}({\cal F})$, and the action probability when no features are known, $\tilde{\pi}_{s}^{a}(\emptyset)$, is fully distributed over the features ${\cal F}$:
$\tilde{\pi}_{s}^{a}({\cal F})-\tilde{\pi}_{s}^{a}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\pi}_{s}^{a}).$
Axiom 11 (Linearity)
@@
For a composite policy $\alpha\,\pi+(1-\alpha)\,\omega$ that follows, at each decision stage, policy $\pi$ with probability $\alpha$ and policy $\omega$ with probability $(1-\alpha)$, a state feature’s contribution to the action probability $\alpha\,\pi(s,a)+(1-\alpha)\,\omega(s,a)$ is equal to the weighted sum of its contributions to the individual action probabilities $\pi(s,a)$ and $\omega(s,a)$:
@@
Axiom 12 (Symmetry)
Two features are assigned the same contribution to the action probability $\pi(s,a)$ if they make equal marginal contributions to the probability of selecting action $a$ in state $s$ for all possible coalitions of features:
@@
Axiom 13 (Nullity)
A feature is assigned zero contribution to the action probability $\pi(s,a)$ if it makes zero marginal contribution to the probability of selecting action $a$ in state $s$ for all possible coalitions of features:
@@
These four axioms are directly analogous to the standard axioms from cooperative game theory used to fairly assign player contributions to the outcome of a coalitional game, as discussed in Section 2.2.
@@
We now revisit our autonomous vehicle example introduced in Sections 3 and 1. At the first junction, the vehicle turns right. How can we explain this behaviour? Before we can apply our framework to attribute the decision to the observed features at that state, we must first consider how the agent might behave when one or more features are unknown.
We examine two possible ways the agent could act at the first junction when features are unknown. For clarity, we refer to these behaviours as separate agents. Agent A bases its decisions solely on the direction the sign indicates; if the direction is visible, it follows it. Otherwise, it chooses randomly. Agent B uses both direction and distance to determine which way to turn; if it observes a distance of 10 miles or a direction of right, it moves right. Otherwise, it also chooses randomly. These assumptions give rise to different policies under unknown information, shown in Table 1.
@@
A $\pi(\text{L})=0,\ \pi(\text{R})=1$ 0, 10,\ 10 , 1 0.5, $,\ 0.5$ 0.5 0.5, $,\ 0.5$ 0.5
B 0, 10,\ 10 , 1 0, 10,\ 10 , 1 0, 10,\ 10 , 1 0.5, $,\ 0.5$ 0.5
Table 1: Agent behaviour at the first junction under different subsets of observed features. Each cell shows the probability of turning left (L) and right (R).
We then compute the Shapley values for each policy, attributing the agent’s choice of action to the direction and distance features. The results are shown in Table 2. For Agent A, all credit is assigned to the direction feature since the policy disregards distance entirely. For Agent B, both features contribute equally. In both cases, the features increasing the probability of turning right must decrease the probability of turning left, resulting in negative Shapley values for turning left. This example illustrates how the explanation depends critically on what is assumed about agent behaviour under partial information. We next discuss how we might approach this problem in general.
@@
A L 0.5-0.5- 0.5 0
R $0.5=\tfrac{1}{2}[(1-0.5)+(1-0.5)]$ 0
B L 0.25-0.25- 0.25 0.25-0.25- 0.25
R 0.250.250.25 0.250.250.25
Table 2: Shapley values at the first junction for each feature under the two agent policies.
@@ 4.2 What Can Be Assumed About Agent Behaviour When Some State Features Are Unknown?
We now consider how an agent behaves when some state features are unknown—a question that must be addressed to compute Shapley values. Several approaches might be considered. A natural proposal is to define new learning problems over subsets of observed features—retaining the original environment’s dynamics and rewards—and train policies using reinforcement learning. However, such policies may fail to converge due to partial observability. Even when they do, if the original policy was suboptimal—or not learnt via rewards at all—the newly learned policy may not meaningfully correspond to how the original agent would behave with unknown features, since it solves a different optimisation problem. Another proposal is introducing a special feature value to indicate that a feature is unknown, possibly implemented by replacing unknown features with a fixed placeholder in the agent’s input. But without adapting training, the agent has no reason to interpret such values as indicating missingness. These approaches both require changes to the MDP or training process, limiting generality and making them unsuitable for post hoc explanation—though they may offer promising directions for future work. In what follows, we introduce a general and post hoc approach to evaluating behaviour under unknown feature values.
@@
Definition 17 (Discrete Policy Characteristic Function)
@@
$\tilde{\pi}_{s}^{a}(\mathcal{C})\coloneqq\mu(s_{\mathcal{C}},a)=\mathbb{E}[\pi(S,a)\,|\,S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\pi(s^{\prime},a).$
@@
The expectation is taken over the marginal steady-state distribution $p^{\pi}(S)$: the probability that an agent following policy $\pi$ is in state $s$.
When only the features in $\mathcal{C}$ are known, the policy $\mu$ provides the closest approximation to the original policy $\pi$. As such, the difference between $\mu(s_{\mathcal{C}},a)$ and $\pi(s,a)$ quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are to reproduce the behaviour of policy $\pi$; this reflects the influence of those unknown features. Defining the characteristic function as $\tilde{\pi}_{s}^{a}(\mathcal{C})=\mu(s_{\mathcal{C}},a)$ thus ensures that the differences across feature subsets used to compute Shapley values reflect the influence of the unknown features, making it a principled choice for the discrete policy game. Because this construction relies only on the agent’s policy, not on how it was derived, it is general to any policy—optimal and non-optimal.
Unlike defining new learning problems with missing features, or imputing them with a masking value, representing unknown features by marginalising over their conditional distribution does not require changing the MDP or training process. An equally general but alternative approach to representing unknown feature values is to marginalise over their joint distribution:
@@
This approach differs from marginalising over the conditional steady-state distribution in much the same way as in supervised learning (Section 2.3). It assumes feature independence and yields equivalent results when that assumption holds. As a result, it assigns zero Shapley value to any feature not explicitly used by the policy. This outcome may appear desirable, but in most function approximation settings, it is less meaningful: such models tend to incorporate all available features to some degree, making true feature irrelevance rare except in contrived examples. When features are dependent, joint marginalisation may produce combinations of values that do not correspond to any valid state. Shapley value attribution relies on measuring behavioural changes in response to changes in known feature values; however, querying a policy with an undefined state renders these changes uninterpretable. For this reason, we focus on the conditional approach in this work, although adapting our framework for joint marginalisation is straightforward. The choice of characteristic function may ultimately depend on the user, but to interpret the resulting Shapley values correctly, it is of paramount importance that they understand the underlying assumptions.